# Algorithm Generation via Creative Ideation

### v1

@@ 3.3 Developing Stimuli into Executable Solutions
Third, developing executable solutions from stimuli involves structured checkpoint-based steps, or waypoint reasoning. Waypoints are intermediate questions serving as checkpoints, which LLMs need to answer sequentially along the reasoning path. We observe that waypoints prevent LLMs from superficially developing solutions, e.g., simply turning stimuli into variable names in code. This addresses the problem with simply using free-form chain-of-thought (Wei et al., 2023; Mehrotra et al., 2024).
Our current implementation specifies the following waypoints. The first waypoint focuses on property extraction, where LLMs are prompted to reason about concepts and properties of the given stimuli. Considering the keyword “flower”, GPT-4o outputs “a ring of petals” as one property. The second waypoint focuses on problem mapping, and LLMs reason about problem-related observations that can be derived from these concepts and properties. In the previous example, GPT-4o outputs an observation that a cache can use a circular buffer. The third waypoint is solution formulation, which prompts LLMs to generate the complete description of a new solution. The observation derived from “flower” would be combined with other observations, for a complete solution. The final waypoint is code generation, and LLMs turn the solution description into executable Python code.
@@ 4 Empirical Results
Baselines.
We have 21 extensive baselines. LLM-based heuristics design baselines include MCTS-AHD (Zheng et al., 2025), ReEvo (Ye et al., 2024), OpenEvolve (Sharma, 2025), PlanSearch (Wang et al., 2025), and Repeated Sampling (Brown et al., 2024). We also include human-crafted SOTA heuristics for cache replacement problem (LRU, LFU, FIFO, Sieve (Zhang et al., 2024), S3FIFO (Yang et al., 2023), TinyLFU (Einziger et al., 2017), SLRU (Huang et al., 2013), Clock (Corbato, 1969), and ARC (Megiddo et al., 2003)), and online bin packing problem (Next Fit, Worst Fit, Almost Worst Fit (AWF) (Johnson, 1973), First Fit (Dósa & Sgall, 2013), Best Fit (Dósa & Sgall, 2014), Harmonic-k (Lee & Lee, 1985), and Refined First Fit (RFF) (Yao, 1980)). For Harmonic-k, we set =4 to align its number of bin categories with RFF.
MetaMuse.
We take a dictionary of common English words and remove stop-words to get 2,899 keywords. Each solution is ideated from =4 keywords. The feedback embedding consists of performance measurements on =30 traces. For cache replacement, these traces are generated by libCacheSim (Yang et al., 2020), from different Zipfian distributions. For bin packing, traces are from various Weibull distributions. The number of RSDict-SF warmup solutions is =100, roughly one-third of the total solutions in one experiment. Finally, prompts are included in §A.
Experiment setup.
We focus on two high-impact problems at a global cloud provider: cache replacement and bin packing. In each experiment, all methods aim to ideate and generate 350 executable solutions. To the best of our ability, we configure PlanSearch to output at least 15, 11, and 12 observations, while ideating with GPT-4o, Llama3.3-70B, and DeepSeek-V3, respectively. Many baselines such as MCTS-AHD can tune parameters in their solutions. §5 discusses safeguards against unsafe solutions, and environments are instrumented to catch errors, e.g., long-running execution. Unsafe solutions are re-implemented. Coding prompts are in §A.2.
To evaluate cache replacement solutions, we use 96 real-world workload traces from four data access scenarios: RetrievalAttention (Liu et al., 2024a) (24 “ra-fwe” and 24 “ra-multikey” traces), Tencent block storage (Zhang et al., 2020; 2018) (24 “tencent-storage” traces), and Alibaba cloud storage (Alibaba, 2020; Li et al., 2020; Wang et al., 2022) (24 “alibaba-storage” traces). The cache capacity is set to 10% of the number of distinct objects in each trace.
@@ 4.1 MetaMuse Generates High-performing Solutions
Cache replacement.
We first delve into results from GPT-4o. At the 90th-percentile trace, MetaMuse achieves 5.17%–9.89% lower miss ratio than LLM-based baselines, and 1.75%–13.03% lower than human heuristics. At the 75th-percentile trace, MetaMuse achieves 3.62%–6.39% lower miss ratio than LLM-based baselines, and 6.62%–35.76% lower than human heuristics. We also see the improvement with Llama3.3-70B — at the 90th-percentile trace, MetaMuse achieves 2.41%–7.20% lower miss ratio than LLM-based baselines, and 0.94%–10.34% lower than human heuristics; at the 75th-percentile trace, MetaMuse achieves 3.67%–5.17% lower miss ratio than LLM-based baselines, and 5.48%–34.62% lower than human heuristics. On DeepSeek-V3, at the 90th-percentile trace, MetaMuse achieves up to 6.05% lower miss ratio than LLM-based baselines, and up to 9.14% lower than human heuristics.
Online bin packing.
Figure 4 compares top solutions, as selected by the average bin usage over all 96 workload traces. Box plots show their bin usage reduction, with respect to First Fit heuristics.
We first delve into results from GPT-4o. At the 90th-percentile trace, MetaMuse achieves 9.25%–9.42% lower bin usage than LLM-based baselines, and 9.25%–20.59% lower than human heuristics. At the 75th-percentile trace, MetaMuse achieves 6.94%–7.75% lower bin usage than LLM-based baselines, and 7.56%–18.36% lower than human heuristics. On Llama3-7B, at the 90th-percentile trace, MetaMuse achieves up to 0.19% less bin usage than LLM-based baselines, and up to 12.50% lower bin usage than human heuristics. On DeepSeek-V3, at the 90th-percentile trace, MetaMuse achieves up to 21.06% less bin usage than LLM-based baselines, and up to 30.93% lower bin usage than human heuristics.
@@ 4.2 MetaMuse Generates the Most Diverse Set of Solutions
Cache replacement.
The discovery of useful solutions depends on having diverse solutions. We evaluate diversity by the number of solutions with a distinct feedback embedding. Across all LLMs, MetaMuse consistently achieves higher diversity over LLM-based baselines. From analyzing 350 solutions generated by each method on GPT-4o, MetaMuse has 1.47 more distinct solutions on average. On DeepSeek-V3, MetaMuse has 1.57 more distinct solutions on average. On Llama3.3-70B, results show that MetaMuse can have 1.78 more distinct solutions on average.
Furthermore, we observe that having a higher diversity translates into lower availability bias. We recall that Repeated Sampling tends to generate solutions that tightly cluster to well-known human heuristics such as LRU, LFU, and FIFO, due to availability bias (§2.3). Using the nine human heuristics as centroids, empirical results show that MetaMuse reduces the cluster density (computed by ), for nearly all centroids. In other words, MetaMuse solutions are less tightly clustered than Repeated Sampling. For example, MetaMuse reduces the LRU cluster density by 35.25% (i.e., 6.15 down to 3.98). We see cluster density reduction even on the less common heuristics, e.g., the ARC cluster density reduces by 77.45% (i.e., 1.02 down to 0.23).
Online bin packing.
We evaluate diversity by the number of solutions with a distinct feedback embedding. Across all LLMs, MetaMuse consistently achieves a higher diversity over LLM-based baselines. From analyzing 350 solutions generated by each method on GPT-4o, MetaMuse has 1.44 more distinct solutions on average. On DeepSeek-V3, it has 1.80 more distinct solutions on average. On Llama3-70B, it has 1.31 more distinct solutions on average.
4.3 MetaMuse Has a Low Per-solution Cost
We see having a low per-solution cost, as a factor driving the adoption of LLMs for creative ideation. We illustrate with the case of generating cache replacement solutions. The total cost for MetaMuse to generate one full solution is up to 4.93 cents with GPT-4o, up to 3.80 cents with DeepSeek-V3, and up to 3.85 cents with Llama3.3-70B. Looking at the cost breakdown, it costs 3.72 cents for GPT4-o to output one implementation in Python. We note that if an implementation is considered unsafe by our safeguards (§5), we re-implement it.
@@ 4.4 Surprising Designs from MetaMuse
In addition to quantitative evaluations, we highlight surprising designs not immediately obvious to engineers. We consider a top-performing cache replacement solution: MetaMuse-533 (§B.1).
One is the NSE counter tracking the number of times that a cached object has witnessed eviction events. It contributes to the score calculation, along with recent access time and frequency counter. Due to NSE, MetaMuse-533 behaves to favor objects that have remained longer in the cache (up to some thresholds). Interestingly, NSE seems to counter the common belief of favoring newly admitted objects, in order to learn their access patterns and prevent thrashing (Denning, 1968).
Another is the use of saturating counters, which increment until some pre-defined threshold. The typical usage is to reduce storage overhead by capping the maximum value of a counter (Yang et al., 2023; Corbato, 1969). Instead, MetaMuse-533 seems to use saturating counters to accumulate meaningful usage history, but not so much as to mislead eviction decisions. For example, a saturating frequency counter allows MetaMuse-533 to identify bursty objects, while preventing their counts from growing too large during the burst.
@@ 4.5 Ablation Study
We test key components in MetaMuse, with the case study of cache replacement algorithms.
Stimuli selection strategies.
We test RSDict and RSDict-SF by running experiments on GPT-4o. First, the majority of RSDict and RSDict-SF solutions outperform Repeated Sampling solutions, and this demonstrates the value of having stimuli for steering ideation. Comparing solutions for the 90th-percentile workload trace, 67.20% of RSDict solutions and 75.60% of RSDict-SF solutions have a higher cache hit ratio than Repeated Sampling. In addition, these hit ratios can be up to 13.13% and 27.07% higher, respectively. Similarly, for the 75th-percentile workload trace, we see 66.80% of RSDict solutions and 72.40% of RSDict-SF solutions outperform Repeated Sampling. In addition, these hit ratios can be up to 15.43% and 21.93% higher, respectively.
Second, compared to RSDict, RSDict-SF has a greater number of solutions that can outperform Repeated Sampling. Empirical results show 12.50% and 8.38% more solutions, with respect to the 90th percentile and 75th percentile of workload traces, respectively. Furthermore, looking at solution diversity, RSDict-SF results in 13.17% more distinct solutions than RSDict.
Waypoint reasoning.
Figure 5 shows that waypoint reasoning enables MetaMuse to generate more diverse solutions. We fix the coding agents to GPT-4o. At each iteration, we instrument MetaMuse to generate solutions with and without waypoint reasoning. When ideating with GPT-4o (5(a)), waypoint reasoning improves the number of distinct solutions from 149 to 175 for RSDict, and from 152 to 197 for RSDict-SF. When ideating with DeepSeek-V3 (5(c)), waypoint reasoning improves the number of distinct solutions from 113 to 144 for RSDict, and from 119 to 140 for RSDict-SF. When ideating with Llama3.3-70B (5(b)), waypoint reasoning improves the number of distinct solutions from 124 to 148 for RSDict, and from 129 to 154 for RSDict-SF.
@@ 5 Discussion
Safeguards against unsafe solutions.
In the case of system algorithms, three main errors result in unsafe implementations: (1) long-running execution (e.g., more than 5 seconds), (2) excessive memory usage, and (3) illegal behaviors, such as falsely claiming a cache hit when the requested object is not present. To this end, our environments are configured to monitor the process execution time and memory usage. Furthermore, we constrain the coding scope for LLMs by explicitly specifying the functions (e.g., insert and evict for caching algorithms). Finally, with Python scripts as outputs, we can leverage testing and verification tools in the Python ecosystem.
Limitations and future discussion.
First, our implementation explores an instantiation of external stimuli (§3.2), which is specifically chosen for being general and problem-agnostic. However, the MetaMuse framework is extensible to take in other instantiations, and future work will explore possibilities such as long-term task-related memories. Second, our work currently focuses on fully automated ideation, and an interesting direction is to explore whether subjective human guidance (e.g., human expertise) could further improve the solution usefulness. Third, we evaluate on real-world workload traces (§4), and deploying LLM-generated algorithms to a business-critical production environment remains an ongoing collaboration with engineers at a global cloud provider. We look forward to leverage lessons learned for deeper practical impacts.

### v2

@@ 3.3 Developing Stimuli into Executable Solutions
Third, developing executable solutions from stimuli requires structured waypoint reasoning. Waypoints serve as intermediate checkpoints that progressively transform seemingly unrelated stimuli (§3.2.1) into solutions for the problem at hand. Unlike free-form chain-of-thought (Wei et al., 2023; Mehrotra et al., 2024), we observe that waypoints prevent LLMs from superficially developing solutions, e.g., simply turning stimuli into variable names in code.
MetaMuse currently defines the following waypoints. The first is property extraction, where the LLM associates the given stimuli with related concepts and properties. For example, given the keyword “flower,” one associated property might be “a ring of petals.” These outputs are then passed to the second waypoint, problem mapping, where the LLM connects the extracted properties to problem-related observations. Continuing the example, “a ring of petals” may remind the LLM of a circular structure, which can be mapped to a circular buffer in algorithm design. These observations are then fed to the third waypoint, solution formulation, where the LLM combines observations and synthesizes the complete description of a new solution. The final waypoint is code generation, where the LLM translates the solution description into executable Python code. The waypoint prompts are included in the appendix (§C.2).
@@ 4 Empirical Results
Baselines. We evaluate MetaMuse against 21 baselines. The LLM-based heuristic design baselines include MCTS-AHD (Zheng et al., 2025), ReEvo (Ye et al., 2024), OpenEvolve (Sharma, 2025), PlanSearch (Wang et al., 2025), and Repeated Sampling (Brown et al., 2024). We also include human-crafted state-of-the-art heuristics for cache replacement, including LRU, LFU, FIFO, Sieve (Zhang et al., 2024), S3FIFO (Yang et al., 2023), TinyLFU (Einziger et al., 2017), SLRU (Huang et al., 2013), Clock (Corbato, 1968), and ARC (Megiddo et al., 2003). For the online bin packing problem, we include Next Fit, Worst Fit, Almost Worst Fit (AWF) (Johnson, 1973), First Fit (Dósa & Sgall, 2013), Best Fit (Dósa & Sgall, 2014), Harmonic-k (Lee & Lee, 1985), and Refined First Fit (RFF) (Yao, 1980). For Harmonic-k, we set =4 to align its number of bin categories with RFF.
MetaMuse. We take a dictionary of common English words and remove stop-words to get 2,899 keywords. Each solution is ideated from =4 keywords. The feedback embedding consists of performance measurements on =30 traces. For cache replacement, these traces are generated by libCacheSim (Yang et al., 2020), from different Zipfian distributions. For bin packing, traces are from various Weibull distributions. The number of RSDict-SF warmup solutions is =100, roughly one-third of the total solutions in one experiment. Finally, prompts are included in §C.
Experiment setup. We focus on two high-impact problems at a global cloud provider: cache replacement and bin packing. In each experiment, all methods aim to ideate and generate 350 executable solutions. To the best of our ability, we configure PlanSearch to output at least 15, 11, and 12 observations, while ideating with GPT-4o, Llama3.3-70B, and DeepSeek-V3, respectively. Many baselines such as MCTS-AHD can tune parameters in their solutions. §5 discusses safeguards against unsafe solutions, and environments are instrumented to catch errors, e.g., long-running execution. Unsafe solutions are re-implemented.
To evaluate cache replacement solutions, we use 96 real-world workload traces from four data access scenarios (Table 1): RetrievalAttention (Liu et al., 2024a) (24 “ra-fwe” and 24 “ra-multikey” traces), Tencent block storage (Zhang et al., 2020; 2018) (24 “tencent-storage” traces), and Alibaba cloud storage (ali, 2022; Li et al., 2020; Wang et al., 2022) (24 “alibaba-storage” traces). The cache capacity is set to 10% of the number of distinct objects in each trace.
@@ 4.1 MetaMuse Generates High-performing Solutions
Cache replacement. Figure 3 compares top solutions, as selected by the average cache miss ratio over all 96 workload traces. Box plots show their miss ratio reduction with respect to FIFO.
We first delve into results from GPT-4o. At the 90th-percentile trace, MetaMuse achieves 5.17%–9.89% lower miss ratio than LLM-based baselines, and 1.75%–13.03% lower than human heuristics. At the 75th-percentile trace, MetaMuse achieves 3.62%–6.39% lower miss ratio than LLM-based baselines, and 6.62%–35.76% lower than human heuristics. We also see improvement with Llama3.3-70B—at the 90th-percentile trace, MetaMuse achieves 2.41%–7.20% lower miss ratio than LLM-based baselines, and 0.94%–10.34% lower than human heuristics; at the 75th-percentile trace, MetaMuse achieves 3.67%–5.17% lower miss ratio than LLM-based baselines, and 5.48%–34.62% lower than human heuristics. On DeepSeek-V3, at the 90th-percentile trace, MetaMuse achieves up to 6.05% lower miss ratio than LLM-based baselines, and up to 9.14% lower than human heuristics.
Online bin packing. Figure 4 compares top solutions, as selected by the average bin usage over all 96 workload traces. Box plots show their bin usage reduction, with respect to First Fit.
We first delve into results from GPT-4o. At the 90th-percentile trace, MetaMuse achieves 9.25%–9.42% lower bin usage than LLM-based baselines, and 9.25%–20.59% lower than human heuristics. At the 75th-percentile trace, MetaMuse achieves 6.94%–7.75% lower bin usage than LLM-based baselines, and 7.56%–18.36% lower than human heuristics. With Llama3-7B, at the 90th-percentile trace, MetaMuse achieves up to 0.19% less bin usage than LLM-based baselines, and up to 12.50% lower bin usage than human heuristics. With DeepSeek-V3, at the 90th-percentile trace, MetaMuse achieves up to 21.06% less bin usage than LLM-based baselines, and up to 30.93% lower bin usage than human heuristics.
@@ 4.2 MetaMuse Generates the Most Diverse Set of Solutions
Cache replacement. The discovery of useful solutions depends on having diverse solutions. We evaluate diversity by the number of solutions with a distinct feedback embedding. Across all LLMs, MetaMuse consistently achieves higher diversity over LLM-based baselines. From analyzing 350 solutions generated by each method on GPT-4o, MetaMuse has 1.47 more distinct solutions on average. On DeepSeek-V3, MetaMuse has 1.57 more distinct solutions on average. On Llama3.3-70B, results show that MetaMuse can have 1.78 more distinct solutions on average.
Furthermore, we observe that having a higher diversity translates into lower availability bias. We recall that Repeated Sampling tends to generate solutions that tightly cluster to well-known human heuristics such as LRU, LFU, and FIFO, due to availability bias (§2.3). Using the nine human heuristics as centroids, empirical results show that MetaMuse reduces the cluster density (computed by ), for nearly all centroids. In other words, MetaMuse solutions are less tightly clustered than Repeated Sampling. For example, MetaMuse reduces the LRU cluster density by 35.28% (i.e., 6.15 down to 3.98). We see cluster density reduction even on the less common heuristics, e.g., the ARC cluster density reduces by 77.45% (i.e., 1.02 down to 0.23).
Finally, Figure 5 uses human heuristics as centroids, and evaluate whether generated solutions tend to cluster to a subset of centroids. Ideally, the cluster sizes should be uniform. The standard deviation of cluster size percentages for MetaMuse is 0.08/0.06/0.06, for GPT-4o/DeepSeek-V3/Llama3-70B. This is much lower than baselines: Repeated Sampling (0.14/0.11/0.11), Repeated Sampling with previous solutions (0.11/0.11/0.11), ReEvo (0.14/0.10/0.10), MCTS-AHD (0.17/0.08/0.08), OpenEvolve (0.19/0.19/0.19), and PlanSearch (0.12/0.10/0.99).
Online bin packing. We evaluate diversity by the number of solutions with a distinct feedback embedding. Across all LLMs, MetaMuse consistently achieves a higher diversity over LLM-based baselines. From analyzing 350 solutions generated by each method on GPT-4o, MetaMuse has 1.44 more distinct solutions on average. On DeepSeek-V3, it has 1.80 more distinct solutions on average. On Llama3-70B, it has 1.31 more distinct solutions on average.
4.3 MetaMuse Incurs a Low Per-solution Cost
We view a low per-solution cost as an important factor for driving the adoption of LLMs for creative ideation tasks. In the cache replacement case study, the average cost for MetaMuse to generate one full solution is 2.16 cents with GPT-4o, 2.11 cents with DeepSeek-V3, and 2.35 cents with Llama3.3-70B. These costs include using GPT-4o for code generation. For reference, running the Repeated Sampling baseline (with previous solutions) costs an average of 3.38 cents per solution.
Specifically, property extraction and problem mapping consume an average of 589.39/141.54 input/output tokens per solution with GPT-4o, 982.60/233.75 input/output tokens with DeepSeek-V3, and 768.20/207.50 input/output tokens with Llama3.3-70B. Solution formulation consumes an average of 934.04/230.21 input/output tokens per solution with GPT-4o, 473.01/425.77 input/output tokens with DeepSeek-V3, and 409.53/353.10 input/output tokens with Llama3.3-70B. Code generation consumes an average of 1463.92/1190.39 input/output tokens per solution with GPT-4o, 2223.72/1937.29 input/output tokens with DeepSeek-V3, and 2259.62/1979.89 input/output tokens with Llama3.3-70B. We note that if an implementation is considered unsafe (§5), it is re-implemented.
@@ 4.4 Surprising Designs from MetaMuse
In addition to quantitative evaluations, we highlight MetaMuse-generated designs that are not immediately obvious to engineers. To do so, we look at top-performing cache replacement solution: MetaMuse-533 (§D.1) and MetaMuse-488 (§D.2).
One is the counter named NSE in MetaMuse-533. NSE deviates from the common belief of favoring newly admitted objects. Instead, it tracks the number of times that a cached object has witnessed eviction events. It is used in the score calculation, along with recent access time and frequency counter. Effectively, MetaMuse-533 behaves to favor objects that have remained longer in the cache (up to some thresholds). Our engineers hypothesize that NSE helps to learn objects’ access patterns and prevent thrashing (Denning, 1968).
Another is the use of hashing functions to segment the cache space in MetaMuse-488, which differs from the more common multi-tiered architecture. MetaMuse-488 tries to maintain different groups of cache objects (as identified by their key hashes), and exercises per-group replacement policies in different segments. In contrast, engineers tend to use multi-tiered architecture, as a way to differentiate the importance of cache objects.
Finally, we highlight the use of saturating counters in MetaMuse-533. They increment until reaching a pre-defined threshold. MetaMuse-533 deviates from conventional usage scenarios, which are to reduce storage overhead by capping the maximum value of a counter (Yang et al., 2023; Corbato, 1968). Instead, MetaMuse-533 seems to use saturating counters to accumulate meaningful usage history, but not so much as to mislead eviction decisions. For example, a saturating frequency counter allows MetaMuse-533 to identify bursty objects, while preventing their counts from growing too large during the burst.
@@ 4.5 Ablation Study
We test key components in MetaMuse using the case study of cache replacement algorithms.
Stimuli selection strategies. We test RSDict and RSDict-SF by running experiments on GPT-4o. First, the majority of RSDict and RSDict-SF solutions outperform Repeated Sampling solutions, and this demonstrates the value of having stimuli for steering ideation. Comparing solutions for the 90th-percentile workload trace, 67.20% of RSDict solutions and 75.60% of RSDict-SF solutions have a higher cache hit ratio than Repeated Sampling. In addition, these hit ratios can be up to 13.13% and 27.07% higher, respectively. Similarly, for the 75th-percentile workload trace, we see 66.80% of RSDict solutions and 72.40% of RSDict-SF solutions outperform Repeated Sampling. In addition, these hit ratios can be up to 15.43% and 21.93% higher, respectively.
Second, compared with RSDict, RSDict-SF has a greater number of solutions that can outperform Repeated Sampling. Empirical results show 12.50% and 8.38% more solutions, with respect to the 90th percentile and 75th percentile of workload traces, respectively. Furthermore, looking at solution diversity, RSDict-SF results in 13.17% more distinct solutions than RSDict.
Waypoint reasoning. Figure 6 shows that waypoint reasoning enables MetaMuse to generate more diverse solutions. We fix the coding agents to GPT-4o. At each iteration, we instrument MetaMuse to generate solutions with and without waypoint reasoning. When ideating with GPT-4o (6(a)), waypoint reasoning improves the number of distinct solutions from 149 to 175 for RSDict, and from 152 to 197 for RSDict-SF. When ideating with DeepSeek-V3 (6(c)), waypoint reasoning improves the number of distinct solutions from 113 to 144 for RSDict, and from 119 to 140 for RSDict-SF. When ideating with Llama3.3-70B (6(b)), waypoint reasoning improves the number of distinct solutions from 124 to 148 for RSDict, and from 129 to 154 for RSDict-SF.
@@ 5 Discussion
Safeguards against unsafe solutions. In the case of system algorithms, three main errors can result in unsafe implementations: (1) long-running executions (e.g., exceeding 5 seconds), (2) excessive memory usage beyond user-specified limits, and (3) illegal behaviors (e.g., falsely claiming a cache hit when the requested object is absent). Our benchmarking environments monitor the per-run execution time and peak memory usage when evaluating each solution. Both can be collected from the GNU time command. To detect illegal behaviors, we validate the returned object value against the expected value in the trace. From our experience, 2.28% of solutions generated by MetaMuse are classified as unsafe and subsequently discarded.
Finally, with Python scripts as the output, it is possible to leverage libraries in the Python ecosystem. We discuss two ongoing efforts. One is performing unit tests, with libraries such as unittest (uni, 2025). Another is the design-code equivalence; we are exploring the use of LLMs to generate state machines from design descriptions and code implementations for the equivalence check.
Limitations and future discussion. First, our implementation explores an instantiation of external stimuli (§3.2), which is specifically chosen for its generality and problem-agnostic nature. However, the MetaMuse framework is extensible to take in other instantiations, and future work will explore possibilities such as long-term task-related memories. Second, while our work currently focuses on fully automated ideation, an interesting direction is to explore whether subjective human guidance (e.g., human expertise) could further improve the solution usefulness. Third, although we evaluate on real-world workload traces (§4), deploying LLM-generated algorithms in business-critical production environments remains an ongoing collaboration with engineers at a global cloud provider. We anticipate that lessons from this deployment effort will inform deeper practical impact.
Algorithm benchmarking costs. The RSDict-SF strategy considers the feedback embeddings of previous solutions (§3.2.1), which requires benchmarking each solution on workload traces. Such benchmarking can take a significant amount of time and resources, especially if they involve running an end-to-end system with long workload traces. If costs are prohibitively high, RSDict-SF may impact the practicality of MetaMuse. We note, however, that reducing benchmarking overhead is an active area of research in the systems community Akram & Sawalha (2019), with techniques such as discrete-event simulations, hardware acceleration, and parallelization. At the same time, we are exploring stimuli selection strategies that reduce reliance on extensive benchmarking.
Synergy with existing approaches. Although MetaMuse emphasizes creative leaps in the solution space, we recognize the value in related approaches that iteratively refine algorithm designs (Liu et al., 2024b; Ye et al., 2024; Zheng et al., 2025; Romera-Paredes et al., 2024; Novikov et al., 2025; Sharma, 2025). The two paradigms are complementary, in a way similar to human’s divergent and convergent thinking processes. MetaMuse can provide the initial population of solution candidates, which can then be further improved by the iterative refinement approach.